OpenStudy (anonymous):

An automobile traveling along a straight road increases its speed from 79 ft/s to 98 ft/s in 180 ft. If the acceleration is constant, how much time elapses while the auto moves the 180 ft? Answer in units of s

OpenStudy (anonymous):

here is the formula though im sorry i couldnt give you more :/ vf^2=v0^2+2ad

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

would it be (98)^2=79^2+180?

OpenStudy (theeric):

Close! The \(180\) is \(180\ [ft]\), which is the distance \(d\). So where you have \(180\), you should have \(2\ a\times 180\). So put that in, and you can solve for \(a\), the acceleration! Then you can use the definition of acceleration, which is \(a=\dfrac{v-v_0}{\Delta t}\) and solve for \(\Delta t\). You'll have the other three values! :)

OpenStudy (anonymous):

98^2-79^2=2*a*180 because \[v^2-v_0^2 = 2ax\] where x is the distance travelled. With this formula we find a=9.342 m/s^2 And 98=79+9.342*t because \[v=v_0+at\] where a is the acceleration and t is the time elapsed. With this formula we find t=2.034 s Hope this helps. Sincerely.

OpenStudy (anonymous):

Sorry about not seeing this sooner guys o.o my teacher explained it yay!

OpenStudy (theeric):

Haha, it's great that your teacher explained it well anyway! Congrats on getting it, anyway!